OpenStudy (anonymous):

Write x2 - 8x + 13 = 0 in the form (x - a)2 = b, where a and b are integers. a.)(x - 4)2 = 3 b.) (x - 3)2 = 2 c.)(x - 2)2 = 1 d.) (x - 1)2 = 4

hartnn (hartnn):

so you need to complete the square , divide the co-efficient of 'x' by 2, what do u get ?

jigglypuff314 (jigglypuff314):

completing the square for example x^2 - 6x + 6 = 0 you know that x^2 - 6x + 9 = (x-3)^2 then what would you need to do to x^2 - 6x + 6 in order to get it like x^2 - 6x + 9 to factor it? x^2 - 6x + 6 + 3 - 3 right? :P then x^2 - 6x + 9 - 3 = 0 so (x - 3)^2 - 3 = 0 so (x - 3)^2 = 3 sorry for the messiness :P but that's the idea.

OpenStudy (anonymous):

so it's 3?

jigglypuff314 (jigglypuff314):

huh?

hartnn (hartnn):

what 3 ?

OpenStudy (anonymous):

lol meant b*

jigglypuff314 (jigglypuff314):

not quite, I was giving an example. look at how I did it and try to apply it to your question :P so how to you get from x^2 - 8x + 13 to look like x^2 - 8x + 16 ?

hartnn (hartnn):

did you solve it and got b =3 ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

not doing it right i think

hartnn (hartnn):

how did you solve it ? because b= 3 is correct :)

OpenStudy (anonymous):

wouldnt it be A that's what im thinking

hartnn (hartnn):

yes, its A. your thought iscorrect :)

OpenStudy (anonymous):

thanks :)

hartnn (hartnn):

welcome ^_^